# Eccentric Mergers in AGN Discs: Influence of the Supermassive Black-Hole on Three-body Interactions

### v1

@@ 2.2 Effects from tidal fields
where $|E_{0}|$ and $|E_{b}|$ here denote the initial energy of the three-body system and the energy of the IMS binary, respectively. Relating this to the Hill sphere boundary, we now note that the interaction terminates if $a_{bs}>R_{H}$. To calculate the correction from the Hill sphere, we therefore have to calculate the probability that the 3-body system splits into a IMS with $a_{bs}>R_{H}$. For this, we first use that the total initial energy of the three-body system in the considered hard binary limit, is about the binding energy of the initial binary, i.e. $|E_{0}|\approx Gm^{2}/(2a_{0})$. In this limit the energy of a given IMS binary relates to the binary+single SMA $a_{bs}$ by
@@
With this formalism we can now estimate the number of IMS, $\mathcal{N}$, with an without a tidal field. In the absence of the tidal field, $\mathcal{N}$ is given by the ratio between the bound-state area, and the escaper-area,
@@
where we have used the distribution from Eq. 18. In the case of system constrained by a Hill sphere $R_{\rm H}$, the corresponding number of IMS, $\mathcal{N}_{H}$, is instead given by the the ratio:
@@
If we ask at what $R_{\rm H}/a_{0}$ the probability has decreased by a factor of $2$, one finds that happens when $R_{\rm H}/a_{0}\sim 7$. Therefore, if the binary is within a few times the Hill sphere to start out with, then the outcomes from isolated interactions without the tidal perturber are expected to return probabilities that are accurate within a factor of order of unity. With a solid theoretical understanding and models to test, in the section below we treat this problem with full PN N-body simulations.
@@ 3 Model and Methods
Having described and considered various theoretical aspects of how tides might impact the chaotic interaction of binary-single scatterings and their outcomes, we now turn to exploring this problem in more detail using full PN N-body simulations. As the key interest is related to how eccentric mergers are produced at high rates in disc-like environments, where the population of co-planar objects is significant, we start here by focusing on the co-planar setup. After this we move on to out-of-plane configurations in Sec. 4.5. In the following sections we present our adopted model, numerical methods, and results from full N-body scatterings including a nearby SMBH.
@@ 3.1 Initial configuration of binary-single system orbiting the SMBH
  • Setup I (blue): The single BH is initially placed at the outer boundary of the Hill sphere w.r.t the SMBH, so that its distance from the SMBH is $R_{\rm CM}+R_{\rm H}$.
@@ 3.2 Merger classification
The end-state of few-body scatterings can generally be classified and divided into a range of distinct outcomes depending on what environments the scatterings are taking place in, on the probed observables, and on how finite sizes as well as energy dissipative terms (GW radiation, tides, gas, etc.) are included (e.g. Samsing et al., 2017). In this work we only report results for BBH mergers that have a notable eccentricity, $e$, at frequency $f_{\rm gw}$ as these carry unique information about the nature of the formation channel. In addition, when focusing on eccentric BBH mergers in the LVK bands, the merger time is so short that we do not have to complicate the study by introducing and working with several outcome types. For example, if we considered eccentric LISA sources instead, the relevant population would be the BBHs that do not undergo a 3-body merger but instead are left unbound after the interaction with a notable eccentricity (Samsing & D’Orazio, 2018; D’Orazio & Samsing, 2018). However, in the AGN disc, tracking this population is very difficult, as the evolution of such binaries after the interaction is likely to either be greatly affected by the gas, or by other objects embedded in the disc. In stellar clusters this is in contrast relatively easy as was demonstrated in Samsing (2018). Focusing on eccentric LIGO/Virgo/KAGRA sources in the adopted AGN setup is more robust, as they form and merge promptly. For these reasons, in this paper we mainly focus on the prompt BBH mergers that form in our simulations and appear with notable eccentricity (>0.1) in the LVK bands ($\sim 10\ \rm Hz$).
@@ 3.3 Numerical methods and PN corrections
We adopt the same extensively tested N-body code as in Samsing et al. (2014, 2017); Samsing (2018); Samsing et al. (2022), which includes GR effects by the use of the PN formalism (Blanchet, 2006, 2014). In this formalism the effects from GR are added to the Newtonian acceleration as an expansion in $v/c$,
@@
where bold symbols indicate vectors. Using the same notation as in Blanchet (2006), the separation vector is defined here as $\boldsymbol{r}_{12}=\boldsymbol{r}_{1}-\boldsymbol{r}_{2}$ with corresponding relative velocity vector $\boldsymbol{v}_{12}$, and unit vector $\hat{\boldsymbol{r}}_{12}$ given by $\boldsymbol{r}_{12}/r_{12}$. We note that PN terms have been derived up to higher orders; however, it is not clear how the series converges, and for stability we restrict ourselves to not include higher order terms.
@@ 4 Results
We here present the main results on the formation of eccentric BBH mergers forming as a consequence of binary-single interactions taking place near an SMBH in a disc-like setup. Below we start by showing a few illustrative cases where the tidal field from the SMBH clearly plays a role in the interaction and outcome. We then show the probability for forming eccentric mergers as a function of the BBH semi-major axis (SMA), relative to the Hill sphere of the SMBH, for both co-rotating and counter-rotating interactions. These results are compared to the scatterings by S22.  [^1] We note that in Trani et al. (2023), a large fraction of eccentric mergers were also found in their setup with a SMBH, but the ICs and underlying physical setup (inclusion of stellar-disc velocity dispersion) are different from ours, which makes it difficult to compare side-by-side.. The new features we find are then explored by considering the phase-space, or topology, of the scatterings. Finally, we show results for out-of-plane scatterings, as the binary-single orbital inclination are believed to be critical in the production of eccentric mergers as shown in S22.
@@
Fig. 4 shows two different examples of three-body scatterings for Setup I, where the single BH is approaching the BBH from the outside, as earlier described in Fig. 3. Both scatterings are plotted with respect to the binary+single CM. The top panel displays an interaction that is terminated before its natural outcome as a result of the tidal field from the SMBH, which breaks up the binary as the interaction here leads to an IMS that extents outside of the Hill sphere (blue circle). For this scattering, the initial binary semi-major axis is 1 AU and the distance from the SMBH is chosen such that the size of the Hill radius is 10 AU. The initial binary phase angle $\phi_{0}$ is ${}^{\circ}$. For the experiment in the bottom panel, the SMA and Hill radius are kept the same as in the left, but the binary angle is instead ${}^{\circ}$. As shown, this variation in the binary phase leads to a completely different outcome, namely the merger between two of the three objects (pink, dark-blue) while the single BH is still bound, i.e. the outcome we denote a ‘3-body merger’. This outcome is not possible in a purely Newtonian code, as the inspiral of the two objects undergoing the merger is entirely driven by the radiation arising from the 2.5PN term. Comparing the two shown scatterings, we clearly see that some interactions almost promptly undergo a merger (bottom panel), whereas some enter the resonating state as we saw in Fig. 1. As described in Sec. 2.2, when the system enters these chaotic states the tidal field form the SMBH is likely to split it apart before e.g. a 3-body merger is formed. If gas is present one can imagine the drag forces acting continuously of the bottom panel interaction, which would gradually transfer energy out of the entire system. Such processes could essentially protect the system from undergoing a tidal breakup as it would harden over time. In this simple picture the gas in the accretion disc could therefore help increasing the number of 3-body mergers that otherwise could have terminated earlier by the SMBH tidal field.
@@
Figure 5: Probability for forming a GW merger with $e>0.1$ at peak frequency $f_{\rm gw}\geq$ 10 Hz, for co-rotating (red) and counter-rotating (blue) BBHs in setup I, for $m=100M_{\odot}$ triple BH system. The large black dots are the replicated results from S22. The grey black dots represent the correction factor due to the tidal field according to eq. 22. At lower semi-major axis (up to $\approx$ 0.3 AU) the trend of the probability follows approximately a power law, which then breaks after which curve becomes more fluctuating as the size of the SMA approaches the size of $R_{\rm H}$. Further analysis in Sec 4.3.
@@
Generally, we do not evolve each binary until merger, therefore to classify which BBHs that are falling into the category of having an $e>0.1$ at $f_{\rm gw}\geq 10$ Hz, we take the binary output from the N-body code, which essentially corresponds to the end-state binary SMA and eccentricity, and evolve it using Peters (1964) equations. Specifically, first we estimate what pericenter distance, $r_{f}$, corresponds to a given GW peak-frequency, $f_{\rm gw}$, using Eq. 10. After this, we take the binary SMA and eccentricity outputted from the code, here denoted by $a_{\rm out}$ and $e_{\rm out}$, respectively, and derive the constant $c_{\rm out}$,
@@
We then solve for the eccentricity at which the binary reaches a pericenter distance of $r_{f}$, using the the relation,
@@
For the sources that form with a GW peak frequency above the imposed threshold, .e.g above $10$ Hz, we label these as eccentric mergers as well (S22). Theoretically, these mergers belong to the category of being eccentric in the observable band, whereas from an evolutionary point of view, they could have very different kinds of GW waveforms. For example, mergers with extremely high eccentricity show up as burst sources with a GW signal similar to e.g. GW190521 (Abbott et al., 2020a), whereas less eccentric mergers gradually inspiral throughout the observable band. For the results in Fig. 5, we map both co-rotating (red) and counter-rotating (blue) binaries from Setup I. As one of the main goals with our analysis is to explore if the findings from S22, stating eccentric mergers are produced at exceptionally high probability in AGN-environments, also holds for more realistic ICs and with the inclusion of a SMBH, we here also show data from S22 in black. From this we can now conclude that in terms of the overall scaling and dependency on SMA, the results with a SMBH and with the single approaching from the Hill sphere is similar to the results found in S22, which were based on isolated 3-body scatterings (and no Hill’s boundary). This especially implies that the co-planar geometry that likely is facilitated by the AGN-disc, is still highly effective in creating eccentric GW sources, despite of the influence form the nearby SMBH. Looking closer at the scattering results, we do see some unique features in the data when the SMA of the BBH approaches the Hill sphere $R_{\rm H}$, i.e. in this case when $a_{0}$ approaches $10$ AU. In general, the more we increase the semi-major axis, i.e. the binary size becomes comparable to the Hill radius, the more the probability trend becomes less linear and much more fluctuating. This is not due to numerical noise, but a real feature of the scatterings. There are even regions in which the probability increases with increasing binary size (e.g. after 1 AU, where $\rm a_{0}/\rm R_{\rm H}=0.1$), to then start decreasing again (after 3 AU). Although a direct comparison cannot be made, Boekholt et al. (2023) also found clear (fractal) structures in their two-body Jacobi captures, which also results from keeping track and varying the ICs in a systematic way as we are also doing, but here in the three-body problem.
@@ 4.3 Effects of periodic orbital encounters
To explore the nature of the fluctuating scattering results seen in Fig. 5, we first consider Fig. 7, which further expands on the results by showing the probability for an eccentric merger for two different BH masses (15 $M_{\odot}$, top panels and 100 $M_{\odot}$, bottom panels) and for Setup I (left panels) and Setup II (right panels). All four configurations show similar behavior to the one described in Fig. 5, i.e. at lower SMA, the probability mostly follows a linear trend, while it becomes increasingly more fluctuating as we widen the BBH semi-major axis. The general reason is that our setup, where the single approaches the BBH from the Hill sphere, is so constrained that what we are seeing are real features in the underlying phase-space or topology of the three-body system (Samsing et al., 2018a). Before studying the topology of the scatterings, we first explore analytically if we can estimate when the scatterings go from being a near smooth power-law, to greatly varying. For this, we note that the scatterings most likely have some underlying periodic features, as we keep the initial distance between the binary and single fixed ($R_{\rm H}$ does not change with the BBH SMA) while changing the SMA of the target BBH. This setup gives rise to periodic interactions, where the single falls from the Hill sphere to interact with the BBH after the BBH has undergone $n$ half-revolutions as the SMA decreases. In other words, we might expect periodic or repeating features when the following condition is met for integer $n$,
@@
The values of $a_{n}$ for integer values of $n$ is plotted over the probability distributions shown in Fig. 7 with black vertical solid lines. The vertical lines line up well across the different combinations of ICs, with the fluctuating and repeating features in the probability curves. Furthermore, because of the scaling from Eq. 31, the BBH takes more and more turns, and the spacing between the lines becomes smaller and smaller, with decreasing SMA, which explains why in this limit the probability curves have less features and more closely follows a simple power-law. The more BBH turns and higher line density for lower binary semi-major axis implies that the system increasingly loses information about the exact ICs, in which case it approaches the more chaotic limit that, e.g., was shown in S22 to follow a smoother curve. Correspondingly, when the SMA increases towards $R_{\rm H}$, the impact from the exact ICs are important and the underlying phase-space features start to show up as features in the probability curves. These observations and the theoretical model from Sec. 2.2 are all complimentary, and clearly indicate that, at least for this case, the largest visible effect in the probability when introducing the SMBH is not ‘tidal breakup’, but instead fluctuations from the underlying phase-space. However, it might be that the real astrophysical environment gives rise to more chaotic ICs, e.g. if the single BH does not slowly approach the BBH from the edge of the Hill sphere, in which case the effect from tidal breakup could more the dominant effect. Continuing to consider results from the scatterings shown in Fig. 7, we find that the critical $n$ for which the system can be considered chaotic is for $n\sim 10$.

### v2

@@ 2.2 Effects from tidal fields
where $|E_{0}|$ and $|E_{b}|$ here denote the initial energy of the three-body system and the energy of the IMS binary, respectively. Relating this to the Hill sphere boundary, we now note that the interaction terminates if $a_{bs}>R_{H}$. To calculate the correction from the Hill sphere, we therefore have to calculate the probability that the 3-body system splits into an IMS with $a_{bs}>R_{H}$. For this, we first use that the total initial energy of the three-body system in the considered hard binary limit, is about the binding energy of the initial binary, i.e. $|E_{0}|\approx Gm^{2}/(2a_{0})$. In this limit the energy of a given IMS binary relates to the binary+single SMA $a_{bs}$ by
@@
With this formalism we can now estimate the number of IMS, $\mathcal{N}$, with and without a tidal field. In the absence of the tidal field, $\mathcal{N}$ is given by the ratio between the bound-state area, and the escaper-area,
@@
where we have used the distribution from Eq. 18. In the case of system constrained by a Hill sphere $R_{\rm H}$, the corresponding number of IMS, $\mathcal{N}_{H}$, is instead given by the ratio:
@@
If we ask at what $R_{\rm H}/a_{0}$ the probability has decreased by a factor of $2$, one finds that happens when $R_{\rm H}/a_{0}\sim 7$. Therefore, if the binary is within a few times the Hill sphere to start out with, then the outcomes from isolated interactions without the tidal perturber are expected to return probabilities that are accurate within a factor of order of unity. With a solid theoretical understanding and models to test, in the section below we treat this problem with PN N-body simulations.
@@ 3 Model and Methods
Having described and considered various theoretical aspects of how tides might impact the chaotic interaction of binary-single scatterings and their outcomes, we now turn to exploring this problem in more detail using PN N-body simulations. As the key interest is related to how eccentric mergers are produced at high rates in disc-like environments, where the population of co-planar objects is significant, we start here by focusing on the co-planar setup. After this we move on to out-of-plane configurations in Sec. 4.5. In the following sections we present our adopted model, numerical methods, and results from N-body scatterings including a nearby SMBH.
@@ 3.1 Initial configuration of binary-single system orbiting the SMBH
  • Setup I (blue): The single BH is initially placed at the outer boundary of the Hill sphere w.r.t. the SMBH, so that its distance from the SMBH is $R_{\rm CM}+R_{\rm H}$.
@@ 3.2 Merger classification
The end-state of few-body scatterings can generally be classified and divided into a range of distinct outcomes depending on what environments the scatterings are taking place in, on the probed observables, and on how finite sizes as well as energy dissipative terms (GW radiation, tides, gas, etc.) are included (e.g. Samsing et al., 2017). In this work we only report results for BBH mergers that have a notable eccentricity, $e$, at frequency $f_{\rm gw}$ as these carry unique information about the nature of the formation channel. In addition, when focusing on eccentric BBH mergers in the LVK bands, the merger time is so short that we do not have to complicate the study by introducing and working with several outcome types. For example, if we considered eccentric LISA sources instead, the relevant population would be the BBHs that do not undergo a 3-body merger but instead are left unbound after the interaction with a notable eccentricity (Samsing & D’Orazio, 2018; D’Orazio & Samsing, 2018a). However, in the AGN disc, tracking this population is very difficult, as the evolution of such binaries after the interaction is likely to either be greatly affected by the gas, or by other objects embedded in the disc. In stellar clusters this is in contrast relatively easy as was demonstrated in Samsing (2018). Focusing on eccentric LIGO/Virgo/KAGRA sources in the adopted AGN setup is more robust, as they form and merge promptly. For these reasons, in this paper we mainly focus on the prompt BBH mergers that form in our simulations and appear with notable eccentricity (>0.1) in the LVK bands ($\sim 10\ \rm Hz$).
@@ 3.3 Numerical methods and PN corrections
We perform simulations with the same extensively tested N-body code as in Samsing et al. (2014, 2017); Samsing (2018); Samsing et al. (2022), which adopts a solver based on LSODE with a relative and absolute error set to $10^{-12}$ and an adaptive time-step with a scale based on the relative position, velocity and acceleration of the interacting bodies following standard procedures. The code includes GR effects by the use of the PN formalism (Blanchet, 2006, 2014). In this formalism the effects from GR are added to the Newtonian acceleration as an expansion in $v/c$,
@@
where bold symbols indicate vectors. Using the same notation as in Blanchet (2006), the separation vector is defined here as $\boldsymbol{r}_{12}=\boldsymbol{r}_{1}-\boldsymbol{r}_{2}$ with corresponding relative velocity vector $\boldsymbol{v}_{12}$, and unit vector $\hat{\boldsymbol{r}}_{12}$ given by $\boldsymbol{r}_{12}/r_{12}$. We note that PN terms have been derived up to higher orders; however, it is not clear how the series converges, and for stability we restrict ourselves not to include higher order terms.
@@ 4 Results
We here present the main results on the formation of eccentric BBH mergers forming as a consequence of binary-single interactions taking place near an SMBH in a disc-like setup. Below we start by showing a few illustrative cases where the tidal field from the SMBH clearly plays a role in the interaction and outcome. We then show the probability for forming eccentric mergers as a function of the BBH semi-major axis (SMA), relative to the Hill sphere of the SMBH, for both co-rotating and counter-rotating interactions. These results are compared to the scatterings by S22.  [^1] We note that in Trani et al. (2023), a large fraction of eccentric mergers were also found in their setup with a SMBH, but the ICs and underlying physical setup (inclusion of stellar-disc velocity dispersion) are different from ours, which makes it difficult to compare side-by-side.. The new features we find are then explored by considering the phase-space, or topology, of the scatterings. Finally, we show results for out-of-plane scatterings, as the binary-single orbital inclination is believed to be critical in the production of eccentric mergers as shown in S22.
@@
Fig. 4 shows two different examples of three-body scatterings for Setup I, where the single BH is approaching the BBH from the outside, as earlier described in Fig. 3. Both scatterings are plotted with respect to the binary+single CM. The top panel displays an interaction that is terminated before its natural outcome as a result of the tidal field from the SMBH, which breaks up the binary as the interaction here leads to an IMS that extents outside of the Hill sphere (blue circle). For this scattering, the initial binary semi-major axis is 1 AU and the distance from the SMBH is chosen such that the size of the Hill radius is 10 AU. The initial binary phase angle $\phi_{0}$ is 0. For the experiment in the bottom panel, the SMA and Hill radius are kept the same as in the left, but the binary angle is instead 135. As shown, this variation in the binary phase leads to a completely different outcome, namely the merger between two of the three objects (pink, dark-blue) while the single BH is still bound, i.e. the outcome we denote a ‘3-body merger’. This outcome is not possible in a purely Newtonian code, as the inspiral of the two objects undergoing the merger is entirely driven by the radiation arising from the 2.5PN term. Comparing the two shown scatterings, we clearly see that some interactions almost promptly undergo a merger (bottom panel), whereas some enter the resonating state as we saw in Fig. 1. As described in Sec. 2.2, when the system enters these chaotic states the tidal field from the SMBH is likely to split it apart before e.g. a 3-body merger is formed. If gas is present one can imagine the drag forces acting continuously on the bottom panel interaction, which would gradually transfer energy out of the entire system. Such processes could essentially protect the system from undergoing a tidal breakup as it would harden over time. In this simple picture, the gas in the accretion disc could therefore help increase the number of 3-body mergers that otherwise could have terminated earlier by the SMBH tidal field.
@@ 4.2 Probability of eccentric black-hole mergers
Figure 5: Probability for forming a GW merger with $e>0.1$ at peak frequency $f_{\rm gw}\geq$ 10 Hz, for co-rotating (red) and counter-rotating (blue) BBHs in setup I, for $m=100M_{\odot}$ triple BH system. The large black dots are the replicated results from S22. The grey black dots represent the correction factor due to the tidal field according to eq. 22. At lower semi-major axis (up to $\approx$ 0.3 AU) the trend of the probability follows approximately a power law, that then breaks after which the curve becomes more fluctuating as the size of the SMA approaches the size of $R_{\rm H}$. Further analysis in Sec 4.3.
@@
Generally, we do not evolve each binary until merger, therefore to classify which BBHs are falling into the category of having an $e>0.1$ at $f_{\rm gw}\geq 10$ Hz, we take the binary output from the N-body code, which essentially corresponds to the end-state binary SMA and eccentricity, and evolve it using Peters (1964) equations. Specifically, first we estimate what pericenter distance, $r_{f}$, corresponds to a given GW peak-frequency, $f_{\rm gw}$, using Eq. 10. After this, we take the binary SMA and eccentricity outputted from the code, here denoted by $a_{\rm out}$ and $e_{\rm out}$, respectively, and derive the constant $c_{\rm out}$,
@@
We then solve for the eccentricity at which the binary reaches a pericenter distance of $r_{f}$, using the relation,
@@
For the sources that form with a GW peak frequency above the imposed threshold, e.g. above $10$ Hz, we label these as eccentric mergers as well (S22). Theoretically, these mergers belong to the category of being eccentric in the observable band, whereas from an evolutionary point of view, they could have very different kinds of GW waveforms. For example, mergers with extremely high eccentricity show up as burst sources with a GW signal similar to e.g. GW190521 (Abbott et al., 2020a), whereas less eccentric mergers gradually inspiral throughout the observable band. For the results in Fig. 5, we map both co-rotating (red) and counter-rotating (blue) binaries from Setup I. As one of the main goals with our analysis is to explore if the findings from S22, stating eccentric mergers are produced at exceptionally high probability in AGN-environments, also holds for more realistic ICs and with the inclusion of a SMBH, we here also show data from S22 in black. From this we can now conclude that in terms of the overall scaling and dependency on SMA, the results with a SMBH and with the single approaching from the Hill sphere is similar to the results found in S22, which were based on isolated 3-body scatterings (and no Hill’s boundary). This especially implies that the co-planar geometry that likely is facilitated by the AGN-disc, is still highly effective in creating eccentric GW sources, despite the influence from the nearby SMBH. Looking closer at the scattering results, we do see some unique features in the data when the SMA of the BBH approaches the Hill sphere $R_{\rm H}$, i.e. in this case when $a_{0}$ approaches $10$ AU. In general, the more we increase the semi-major axis, i.e. the binary size becomes comparable to the Hill radius, the more the probability trend becomes less linear and much more fluctuating. This is not due to numerical noise, but a real feature of the scatterings. There are even regions in which the probability increases with increasing binary size (e.g. after 1 AU, where $\rm a_{0}/\rm R_{\rm H}=0.1$), to then start decreasing again (after 3 AU). Although a direct comparison cannot be made, Boekholt et al. (2023) also found clear (fractal) structures in their two-body Jacobi captures, which also results from keeping track and varying the ICs in a systematic way as we are also doing, but here in the three-body problem.
@@ 4.3 Effects of periodic orbital encounters
To explore the nature of the fluctuating scattering results seen in Fig. 5, we first consider Fig. 7, which further expands on the results by showing the probability for an eccentric merger for two different BH masses (15 $M_{\odot}$, top panels and 100 $M_{\odot}$, bottom panels) and for Setup I (left panels) and Setup II (right panels). All four configurations show similar behavior to the one described in Fig. 5, i.e. at lower SMA, the probability mostly follows a linear trend, while it becomes increasingly more fluctuating as we widen the BBH semi-major axis. The general reason is that our setup, where the single approaches the BBH from the Hill sphere, is so constrained that we are seeing real features in the underlying phase-space or topology of the three-body system (Samsing et al., 2018a). Before studying the topology of the scatterings, we first explore analytically if we can estimate when the scatterings go from being a near smooth power-law, to greatly varying. For this, we note that the scatterings most likely have some underlying periodic features, as we keep the initial distance between the binary and single fixed ($R_{\rm H}$ does not change with the BBH SMA) while changing the SMA of the target BBH. This setup gives rise to periodic interactions, where the single falls from the Hill sphere to interact with the BBH after the BBH has undergone $n$ half-revolutions as the SMA decreases. In other words, we might expect periodic or repeating features when the following condition is met for integer $n$,
@@
The values of $a_{n}$ for integer values of $n$ is plotted over the probability distributions shown in Fig. 7 with black vertical solid lines. The vertical lines line up well across the different combinations of ICs, with the fluctuating and repeating features in the probability curves. Furthermore, because of the scaling from Eq. 31, the BBH takes more and more turns, and the spacing between the lines becomes smaller and smaller, with decreasing SMA, which explains why in this limit the probability curves have less features and more closely follows a simple power-law. The more the BBH turns and the higher line density for lower binary semi-major axis imply that the system increasingly loses information about the exact ICs, in which case it approaches the more chaotic limit that, e.g., was shown in S22 to follow a smoother curve. Correspondingly, when the SMA increases towards $R_{\rm H}$, the impact from the exact ICs is important and the underlying phase-space features start to show up as features in the probability curves. These observations and the theoretical model from Sec. 2.2 are all complementary, and clearly indicate that, at least for this case, the largest visible effect in the probability when introducing the SMBH is not ‘tidal breakup’, but instead fluctuations from the underlying phase-space. However, it might be that the real astrophysical environment gives rise to more chaotic ICs, e.g. if the single BH does not slowly approach the BBH from the edge of the Hill sphere, in which case the effect from tidal breakup could be a more dominant effect. Continuing to consider results from the scatterings shown in Fig. 7, we find that the critical $n$ for which the system can be considered chaotic is for $n\sim 10$.